# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v3

@@ A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
Although our evaluation suite pursued different goals, (Abdelaal et al., 2023) served as the primary benchmark against which we compared our work. The authors made significant contributions, establishing a benchmark that set a high standard in the field. However, we identified several critical concerns in their codebase that we aimed to address. Firstly, their comparison did not include hyper-parameter tuning for null imputers, and they did not utilize seeds to control randomization in the null imputers. In contrast, we meticulously tuned hyper-parameters for each null imputer in our study, implemented controlled randomization throughout the entire pipeline using seeds, and rigorously tested the evaluation suite with both unit and integration tests to ensure reproducibility. Secondly, a significant methodological flaw in their benchmark was that null imputers were fitted on the entire dataset without initially splitting it into training and test sets. The dataset split was introduced only during the model training stage, resulting in a clear leakage when null imputers were fitted on samples included in the test set. This issue likely arose because imputers like miss-forest and datawig lack built-in interfaces for fitting only on a training set and then transforming both training and test sets. Presumably, the authors chose to use these imputers without modification. In order to ensure a fair comparison of MVI techniques as they would be used in production systems, in our study, we standardized all 8 MVI techniques to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameter tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable). Below is a summary of the enhancements we implemented for MVI techniques used in our evaluation suite:
@@ B.1. Computing Infrastructure
Our large-scale experimental study incorporated 8 null imputation techniques, 6 model types, 7 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline. This configuration led to an overwhelming total of 23,940 experimental pipelines. The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability. Depending on the dataset size, null imputation technique, and model type, the execution time for a single experimental pipeline ranged from 3 minutes to 16.5 hours. Executing such a large number of pipelines would be impossible without a suitable experimental environment comprising a high-performance computing (HPC) cluster for execution and a 3-node MongoDB cluster for storing experimental results.
Our evaluation suite controller provides a useful separation of experimental pipelines, and the MongoDB cluster gives the ability to store results independently from a local file system. To manage the extensive computational demands, we utilized several optimizations (detailed in Section 3.4) and the SLURM software system, which allowed us to execute up to 100 simultaneous jobs on the HPC cluster. To run experimental pipelines, each SLURM job was allocated 12-24 physical cores or one GPU card, along with 16-96 GB of RAM, depending on the dataset size, imputation method, and model type. For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor.
@@ B.2. Extended Methodology for Simulating Missingness
Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties. Tables 4-10 report the demographic composition of all datasets, specifically, the proportions and base rates of each protected group. Figures 12-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1. For datasets with numerous features, plots include only the most correlated features. These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups. Tables 11-16 detail the missingness scenarios for the folk-employment, folk-income, bank, heart, law-school, and german datasets, respectively.
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
Appendix C Additional Experimental Results for Single-mechanism Missingness
C.1. Additional Accuracy and Fairness Metrics of Predictive Model

### v4

@@ A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
Although our evaluation suite pursued different goals, (Abdelaal et al., 2023) served as the primary benchmark against which we compared our work. The authors made significant contributions, establishing a benchmark that set a high standard in the field. However, we identified several critical concerns in their codebase that we aimed to address. Firstly, their comparison did not include hyper-parameter tuning for null imputers, and they did not utilize seeds to control randomization in the null imputers. In contrast, we meticulously tuned hyper-parameters for each null imputer in our study, implemented controlled randomization throughout the entire pipeline using seeds, and rigorously tested the evaluation suite with both unit and integration tests to ensure reproducibility. Secondly, a significant methodological flaw in their benchmark was that null imputers were fitted on the entire dataset without initially splitting it into training and test sets. The dataset split was introduced only during the model training stage, resulting in a clear leakage when null imputers were fitted on samples included in the test set. This issue likely arose because imputers like miss-forest and datawig lack built-in interfaces for fitting only on a training set and then transforming both training and test sets. Presumably, the authors chose to use these imputers without modification. In order to ensure a fair comparison of MVI techniques as they would be used in production systems, in our study, we standardized our MVI techniques to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameter tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable). Below is a summary of the enhancements we implemented for MVI techniques used in our evaluation suite:
@@
A.3. Datasets and Tasks
bank [^7] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013. It contains information on 40,004 potential customers, with 13 attributes (7 categorical, 6 numerical), including occupation, marital status, education, and a binary target that indicates whether the individual subscribed for a term deposit. Here, age is the sensitive attribute, with $<25$ and $>60$ as the disadvantaged group.
@@ B.1. Computing Infrastructure
Our large-scale experimental study incorporated 15 null imputation techniques, 6 model types, 7 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline. This configuration led to an overwhelming total of 29,736 experimental pipelines. The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability. Depending on the dataset size, null imputation technique, and model type, the execution time for a single experimental pipeline ranged from 3 minutes to 16.5 hours. Executing such a large number of pipelines would be impossible without a suitable experimental environment comprising a high-performance computing (HPC) cluster for execution and a 3-node MongoDB cluster for storing experimental results.
Our evaluation suite controller provides a useful separation of experimental pipelines, and the MongoDB cluster gives the ability to store results independently from a local file system. To manage the extensive computational demands, we utilized several optimizations (detailed in Section 3.4) and the SLURM software system, which allowed us to execute up to 100 simultaneous jobs on the HPC cluster. To run experimental pipelines, each SLURM job was allocated 12-24 physical cores or one GPU card, along with 16-96 GB of RAM, depending on the dataset size, imputation method, and model type. For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor. All evaluated imputation algorithms were implemented using Python 3.9, utilizing their original source code. Some of this code was generously shared directly by the respective authors, for which we express our sincere gratitude. All dependencies are specified in our repository, along with detailed instructions in the README for installing and running the evaluation suite.
@@ B.2. Extended Methodology for Simulating Missingness
Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties. Tables 5-11 report the demographic composition of all datasets, specifically, the proportions and base rates of each protected group. Figures 12-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1. For datasets with numerous features, plots include only the most correlated features. These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups. Tables 12-17 detail the missingness scenarios for the folk-employment, folk-income, bank, heart, law-school, and german datasets, respectively.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix C Additional Experimental Results for Single- and Multi-mechanism Missingness
C.1. Accuracy and Efficiency of MVI Techniques
In this section, we extend Section 4.4 by presenting additional results on the accuracy and efficiency of MVI techniques across single- and multi-mechanism scenarios (S1-S3, S10). Figures 18 and 19 illustrate imputation accuracy, measured by F1 scores for categorical columns and RMSE for numerical columns, respectively. Tables 18–22 summarize the training time of various MVI techniques (in seconds), averaged across these scenarios. In the tables, imputers are sorted by runtime on the folk_emp dataset, while datasets are ordered by their number of rows. Dataset shapes indicate training sets consisting of 70% complete rows and 30% rows with missing values. In the tables, the reported values represent mean runtimes across different seeds, with standard deviations included. Higher F1 scores and lower RMSE and runtimes are desirable for better performance. Notably, boostclean is excluded as it is a joint cleaning-and-training technique that does not generate imputed rows.
Table 18 (same as Table 4 from Section 6, duplicated for convenience) reveals that statistical imputers such as deletion,
median-mode, and median-dummy are the fastest, while still delivering competitive accuracy for larger datasets like heart and folk_emp, as shown in Figures 18 and 19. In contrast, miss-forest, datawig, and auto-ml exhibit the longest training times, with at least one of these methods achieving the highest imputation accuracy in most cases. Interestingly, auto-ml requires three times more training time than datawig, the second most computationally intensive technique. This difference is due to the auto-ML nature of auto-ml, which involves extensive hyperparameter and network architecture tuning.
Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient in terms of runtime. Notably, nomi provides competitive accuracy comparable to miss-forest, datawig, and auto-ml, making it the most optimal technique in terms of the trade-off between imputation accuracy and training time. A particularly noteworthy comparison is between gain and edit-gain. As explained in Appendix A.1, EDIT accelerates the training of parametric imputation models. Applying EDIT to gain results in a 28x speedup on the folk_emp dataset, with even greater improvements for smaller datasets, as shown in Table 18. Importantly, with this runtime acceleration edit-gain achieves accuracy comparable to gain on most datasets, as shown in Figures 18 and 19.
Tables 19–22 provide training times for individual scenarios (S1, S2, S3, and S10). These values are generally consistent with those in Table 18, with minor exceptions: nomi outperforms mnar-pvae and edit-gain in terms of speed for MCAR and MAR scenarios, and clustering is faster than hi-vae in the MAR scenario.
It is worth noting that the training times in Tables 18–22 exhibit substantial standard deviations for many MVI techniques. These deviations arise because the statistics are computed across pipelines executed with different random seeds, resulting in varying training-test splits and model initializations. Consequently, standard deviations can be significant for all MVI techniques. Furthermore, mnar-pvae, not-miwae, miss-forest, datawig, and auto-ml show large standard deviations. For mnar-pvae and not-miwae, their MNAR-specific nature makes them highly sensitive to the data characteristics, leading to variability across splits. In the case of miss-forest, datawig, and auto-ml, the presence of parameters such as maximum iterations or trials introduces early stopping mechanisms that can be influenced by the data from different splits, further contributing to runtime variability.
Table 18: Training time (in seconds) of MVI techniques averaged across single- and multi-mechanism scenarios (S1-3, S10). Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows. Dataset shapes reflect training sets with 30% rows with nulls. Values represent mean runtime across seeds, with standard deviations included.
Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
median-dummy $0.013\pm 0.000$ $0.014\pm 0.001$ $0.021\pm 0.001$ $0.024\pm 0.001$ $0.027\pm 0.000$ $0.053\pm 0.001$ $0.773\pm 0.024$
median-mode $0.012\pm 0.000$ $0.014\pm 0.001$ $0.023\pm 0.001$ $0.025\pm 0.001$ $0.031\pm 0.001$ $0.067\pm 0.003$ $0.933\pm 0.019$
deletion $0.013\pm 0.000$ $0.013\pm 0.001$ $0.024\pm 0.001$ $0.025\pm 0.000$ $0.046\pm 0.001$ $0.087\pm 0.004$ $1\pm 0.049$
mnar__\__pvae $8\pm 0.705$ $14\pm 11$ $14\pm 1$ $22\pm 11$ $55\pm 30$ $42\pm 6$ $206\pm 7$
edit__\__gain $2\pm 0.119$ $2\pm 0.141$ $13\pm 0.221$ $21\pm 2$ $30\pm 1$ $62\pm 7$ $215\pm 4$
nomi $11\pm 3$ $14\pm 7$ $22\pm 2$ $22\pm 2$ $29\pm 1$ $38\pm 2$ $356\pm 20$
tdm $932\pm 74$ $1023\pm 11$ $1297\pm 22$ $1172\pm 92$ $1412\pm 34$ $1310\pm 42$ $1449\pm 31$
notmiwae $161\pm 101$ $217\pm 82$ $555\pm 2$ $804\pm 8$ $665\pm 284$ $1393\pm 535$ $2944\pm 725$
gain $261\pm 12$ $298\pm 5$ $1115\pm 38$ $1484\pm 30$ $1964\pm 36$ $2892\pm 48$ $6148\pm 178$
hivae $68\pm 0.784$ $95\pm 1$ $745\pm 15$ $1073\pm 11$ $2450\pm 130$ $3794\pm 129$ $7163\pm 317$
k__\__means__\__clustering $10\pm 0.402$ $13\pm 0.442$ $27\pm 0.640$ $323\pm 7$ $998\pm 49$ $1037\pm 64$ $7427\pm 778$
miss__\__forest $111\pm 17$ $244\pm 86$ $1758\pm 526$ $2530\pm 851$ $4307\pm 1310$ $6337\pm 1601$ $20358\pm 4934$
datawig $596\pm 185$ $277\pm 59$ $604\pm 46$ $2361\pm 492$ $5089\pm 651$ $7592\pm 854$ $31060\pm 3743$
automl $1953\pm 195$ $1805\pm 212$ $5559\pm 581$ $6803\pm 565$ $13893\pm 1687$ $19055\pm 2710$ $104476\pm 14743$
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 18. F1 of different imputation strategies (colors in the legend) averaged across categorical columns with nulls per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 19. RMSE of different imputation strategies (colors in the legend) averaged across numerical columns with nulls per dataset (x-axis) and missingness mechanisms (subplots). Datasets are ordered in increasing order by size.
Table 19: Training time (in seconds) of MVI techniques for the MCAR scenario (S1). Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows. Dataset shapes reflect training sets with 30% rows with nulls. Values represent mean runtime across seeds, with standard deviations included.
Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
median-dummy $0.015\pm 0.001$ $0.016\pm 0.001$ $0.021\pm 0.000$ $0.025\pm 0.001$ $0.029\pm 0.000$ $0.057\pm 0.002$ $0.846\pm 0.042$
median-mode $0.013\pm 0.001$ $0.016\pm 0.001$ $0.023\pm 0.000$ $0.029\pm 0.001$ $0.034\pm 0.000$ $0.072\pm 0.002$ $1\pm 0.037$
deletion $0.014\pm 0.001$ $0.015\pm 0.000$ $0.027\pm 0.001$ $0.028\pm 0.001$ $0.049\pm 0.001$ $0.100\pm 0.006$ $1\pm 0.036$
nomi $11\pm 4$ $16\pm 9$ $25\pm 2$ $23\pm 2$ $31\pm 1$ $36\pm 2$ $206\pm 5$
edit__\__gain $2\pm 0.133$ $2\pm 0.155$ $14\pm 0.188$ $22\pm 2$ $31\pm 0.792$ $62\pm 7$ $216\pm 4$
mnar__\__pvae $8\pm 1$ $10\pm 1$ $13\pm 1$ $44\pm 43$ $129\pm 121$ $43\pm 0.986$ $228\pm 6$
tdm $897\pm 11$ $1034\pm 6$ $1289\pm 9$ $1095\pm 19$ $1720\pm 68$ $1195\pm 22$ $1448\pm 25$
notmiwae $157\pm 84$ $249\pm 149$ $600\pm 8$ $751\pm 3$ $749\pm 349$ $1472\pm 637$ $2978\pm 578$
gain $257\pm 26$ $261\pm 5$ $1112\pm 72$ $1421\pm 39$ $1902\pm 33$ $2748\pm 48$ $5668\pm 146$
hivae $70\pm 1$ $90\pm 0.228$ $701\pm 8$ $1017\pm 6$ $2531\pm 320$ $3874\pm 86$ $7161\pm 415$
k__\__means__\__clustering $0.921\pm 0.318$ $1\pm 0.326$ $1\pm 0.117$ $18\pm 1$ $26\pm 1$ $114\pm 24$ $7499\pm 720$
miss__\__forest $94\pm 14$ $240\pm 92$ $1837\pm 577$ $2150\pm 989$ $4488\pm 1404$ $6269\pm 1610$ $15154\pm 5008$
datawig $238\pm 69$ $279\pm 66$ $601\pm 70$ $2104\pm 336$ $4654\pm 755$ $7791\pm 921$ $33102\pm 1695$
automl $2219\pm 367$ $1969\pm 131$ $5751\pm 604$ $6936\pm 252$ $14203\pm 1105$ $19315\pm 3505$ $104290\pm 19094$
Table 20: Training time (in seconds) of MVI techniques for the MAR scenario (S2). Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows. Dataset shapes reflect training sets with 30% rows with nulls. Values represent mean runtime across seeds, with standard deviations included.
Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
median-dummy $0.015\pm 0.001$ $0.016\pm 0.000$ $0.021\pm 0.000$ $0.025\pm 0.001$ $0.029\pm 0.000$ $0.057\pm 0.001$ $0.883\pm 0.044$
median-mode $0.013\pm 0.001$ $0.016\pm 0.001$ $0.023\pm 0.000$ $0.027\pm 0.000$ $0.033\pm 0.000$ $0.070\pm 0.003$ $1\pm 0.018$
deletion $0.014\pm 0.001$ $0.015\pm 0.001$ $0.026\pm 0.000$ $0.028\pm 0.000$ $0.048\pm 0.000$ $0.095\pm 0.004$ $1\pm 0.031$
nomi $10\pm 3$ $14\pm 8$ $23\pm 1$ $24\pm 3$ $29\pm 1$ $38\pm 3$ $191\pm 3$
edit__\__gain $2\pm 0.088$ $2\pm 0.120$ $13\pm 0.193$ $20\pm 2$ $30\pm 1$ $63\pm 7$ $209\pm 8$
mnar__\__pvae $8\pm 0.480$ $10\pm 1$ $11\pm 0.646$ $16\pm 0.633$ $35\pm 1$ $41\pm 1$ $213\pm 10$
tdm $907\pm 19$ $984\pm 11$ $1303\pm 8$ $1182\pm 43$ $1335\pm 12$ $1556\pm 57$ $1477\pm 8$
notmiwae $235\pm 140$ $273\pm 0.540$ $586\pm 1$ $797\pm 10$ $803\pm 395$ $1498\pm 553$ $2894\pm 487$
gain $299\pm 11$ $288\pm 4$ $1140\pm 25$ $1590\pm 34$ $1971\pm 34$ $2880\pm 44$ $6659\pm 171$
k__\__means__\__clustering $0.886\pm 0.341$ $1\pm 0.319$ $1\pm 0.122$ $19\pm 2$ $26\pm 1$ $107\pm 22$ $7395\pm 715$
hivae $71\pm 0.788$ $99\pm 1$ $734\pm 13$ $1134\pm 11$ $2337\pm 28$ $3585\pm 84$ $7557\pm 468$
miss__\__forest $112\pm 14$ $251\pm 87$ $1874\pm 606$ $2976\pm 914$ $4493\pm 1218$ $7443\pm 2031$ $19323\pm 5650$
datawig $271\pm 83$ $325\pm 85$ $642\pm 18$ $2398\pm 735$ $4858\pm 695$ $7820\pm 908$ $29815\pm 5383$
automl $2172\pm 187$ $1822\pm 236$ $5495\pm 797$ $7036\pm 507$ $15469\pm 2313$ $19357\pm 3697$ $96490\pm 8667$
Table 21: Training time (in seconds) of MVI techniques for the MNAR scenario (S3). Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows. Dataset shapes reflect training sets with 30% rows with nulls. Values represent mean runtime across seeds, with standard deviations included.
Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
median-dummy $0.015\pm 0.001$ $0.015\pm 0.000$ $0.021\pm 0.001$ $0.025\pm 0.001$ $0.029\pm 0.001$ $0.057\pm 0.002$ $0.773\pm 0.005$
median-mode $0.015\pm 0.000$ $0.016\pm 0.001$ $0.023\pm 0.000$ $0.027\pm 0.000$ $0.034\pm 0.001$ $0.067\pm 0.003$ $0.867\pm 0.006$
deletion $0.014\pm 0.000$ $0.013\pm 0.001$ $0.025\pm 0.000$ $0.027\pm 0.000$ $0.046\pm 0.000$ $0.086\pm 0.004$ $1\pm 0.118$
mnar__\__pvae $8\pm 0.560$ $26\pm 44$ $14\pm 0.825$ $16\pm 0.755$ $34\pm 0.941$ $51\pm 23$ $215\pm 8$
edit__\__gain $2\pm 0.129$ $2\pm 0.144$ $13\pm 0.250$ $21\pm 2$ $30\pm 0.386$ $62\pm 7$ $219\pm 2$
nomi $12\pm 5$ $14\pm 5$ $25\pm 2$ $24\pm 3$ $38\pm 1$ $51\pm 1$ $722\pm 32$
tdm $877\pm 4$ $1050\pm 12$ $1361\pm 51$ $1295\pm 287$ $1312\pm 17$ $1276\pm 73$ $1409\pm 66$
notmiwae $132\pm 100$ $226\pm 96$ $527\pm 1$ $1049\pm 20$ $608\pm 267$ $1518\pm 543$ $3414\pm 814$
gain $248\pm 4$ $349\pm 4$ $1084\pm 29$ $1531\pm 29$ $1875\pm 41$ $2980\pm 56$ $5869\pm 153$
hivae $64\pm 0.228$ $100\pm 1$ $761\pm 7$ $1023\pm 11$ $2529\pm 61$ $3887\pm 82$ $7374\pm 267$
k__\__means__\__clustering $0.887\pm 0.328$ $1\pm 0.323$ $1\pm 0.120$ $18\pm 2$ $26\pm 1$ $106\pm 23$ $7382\pm 805$
miss__\__forest $123\pm 28$ $275\pm 90$ $1442\pm 365$ $2558\pm 818$ $4281\pm 1601$ $4384\pm 972$ $19227\pm 4033$
datawig $247\pm 46$ $232\pm 30$ $581\pm 48$ $2142\pm 380$ $4952\pm 674$ $7048\pm 1119$ $30238\pm 3335$
automl $1704\pm 78$ $1677\pm 309$ $5147\pm 463$ $6707\pm 992$ $13011\pm 2231$ $19299\pm 2315$ $106635\pm 17043$
Table 22: Training time (in seconds) of MVI techniques for the multi-mechanism scenario (S10). Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows. Dataset shapes reflect training sets with 30% rows with nulls. Values represent mean runtime across seeds, with standard deviations included.
Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
median-dummy $0.009\pm 0.000$ $0.010\pm 0.001$ $0.019\pm 0.002$ $0.021\pm 0.000$ $0.021\pm 0.001$ $0.040\pm 0.001$ $0.590\pm 0.007$
median-mode $0.009\pm 0.000$ $0.010\pm 0.000$ $0.023\pm 0.002$ $0.019\pm 0.000$ $0.024\pm 0.000$ $0.059\pm 0.003$ $0.748\pm 0.015$
deletion $0.008\pm 0.000$ $0.009\pm 0.001$ $0.017\pm 0.001$ $0.018\pm 0.000$ $0.039\pm 0.001$ $0.067\pm 0.002$ $0.927\pm 0.013$
mnar__\__pvae $8\pm 0.717$ $9\pm 0.997$ $18\pm 4$ $12\pm 0.505$ $22\pm 0.776$ $33\pm 0.939$ $169\pm 3$
edit__\__gain $2\pm 0.126$ $2\pm 0.145$ $14\pm 0.252$ $21\pm 2$ $29\pm 1$ $62\pm 9$ $215\pm 2$
nomi $10\pm 2$ $13\pm 5$ $17\pm 1$ $16\pm 0.530$ $19\pm 0.962$ $26\pm 1$ $306\pm 38$
tdm $1048\pm 260$ $1024\pm 16$ $1236\pm 21$ $1114\pm 20$ $1281\pm 40$ $1212\pm 16$ $1463\pm 24$
notmiwae $122\pm 81$ $123\pm 83$ $507\pm 0.743$ $618\pm 1$ $502\pm 128$ $1082\pm 406$ $2492\pm 1023$
gain $241\pm 7$ $294\pm 7$ $1126\pm 26$ $1392\pm 19$ $2106\pm 38$ $2960\pm 43$ $6396\pm 241$
hivae $68\pm 0.624$ $93\pm 1$ $785\pm 31$ $1119\pm 16$ $2405\pm 111$ $3828\pm 264$ $6559\pm 119$
k__\__means__\__clustering $39\pm 0.618$ $50\pm 0.800$ $104\pm 2$ $1239\pm 23$ $3914\pm 194$ $3820\pm 185$ $7432\pm 873$
miss__\__forest $114\pm 13$ $207\pm 76$ $1878\pm 553$ $2434\pm 685$ $3966\pm 1017$ $7253\pm 1789$ $27730\pm 5046$
datawig $1627\pm 541$ $272\pm 55$ $591\pm 46$ $2802\pm 518$ $5892\pm 480$ $7708\pm 468$ $31086\pm 4558$
automl $1715\pm 149$ $1750\pm 172$ $5843\pm 461$ $6531\pm 509$ $12888\pm 1099$ $18248\pm 1324$ $110488\pm 14166$
C.2. Additional Accuracy and Fairness Metrics of Predictive Model